OpenStudy (anonymous):

Need Help Please ?? Two different car batteries cost a total of $520. One costs $8 more than twice the other. Which equation could be used to find the cost of each battery? x + 8 = 520 x^2 + x + 8 = 520 x + 2x + 8 = 520 2x + 8 = 520

OpenStudy (anonymous):

nkfnoanfjdsnfjodnsafjoldasnfowlk.dnsfmiajslkfnhsiuoajfkln hauojlfnbdhsiuhofjnkds hfuojnks gay

OpenStudy (lieutenantgeneral):

520 divided by 2 = 260 260 * 8 = 2,080 and/or 520 * 8 = 4,160 2x + 8 = 520